10 volt cell was used to drive a circuit with an external resistor of resistance of 4 ohm and internal resistance of 1 ohm. Determine the current delivered by the cell
1
Expert's answer
2020-06-08T10:27:32-0400

On the figure below "r_i = 1\Omega", "R_e = 4 \Omega".


As far as the resistors are connected in series, the total resistance is:


"R = r_i + R_e = 1+4=5\Omega"

According to the Ohm's law, the current delivered by the cell is:


"I = \dfrac{V}{R} = \dfrac{10V}{5\Omega} = 2A"

Answer. I = 2A.